OpenStudy (anonymous):

how can i apply the difference quotient to this funtion: g(x)=(x)/(x-2) at (3,3)

OpenStudy (anonymous):

u didn't check ur previous post?

OpenStudy (anonymous):

yes i did but, i would like to solve it using this formula: lim as x goes to 0, of (f(x+h)-(f(x))/(h)

OpenStudy (anonymous):

differentiation by definition

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

this is what u want :),

OpenStudy (anonymous):

\[g'={{{x+h} \over{x-2+h}}-{{x} \over{x-2}} \over h}\]

OpenStudy (anonymous):

=(x+h)(x-2)-x(x-2+h)/(x-2+h)(x-2)h

OpenStudy (anonymous):

The top part gets cross multiplied

OpenStudy (anonymous):

by taking the lcm

OpenStudy (anonymous):

x^2+hx-2x-2h-x^2+2x-hx/(x-2+h)(x-2)h =-2h/(x-2+h)(x-2)h =-2/(x-2+h)(x-2) =-2/(x-2)^2 in the loim h ---->0

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Could you help me out one more time

OpenStudy (anonymous):

sure, if i could

OpenStudy (anonymous):

what would be the slope or m

OpenStudy (anonymous):

cause i was trying to find the slope of the functions graph at a given point, namely (3,3)

OpenStudy (anonymous):

slope is the derivative of the function at the given point. just put x=3 in the derivative slope=-2(3-2)^2=-2

OpenStudy (anonymous):

i appreciate your help

OpenStudy (anonymous):

welcome :)